25- for the following systems trasfer functions determines the stability of the system (s2 + 4)2(s+5)
Answer #1

HS) (544)*6+5) (3+4)²=0 s+4 = 0; s+4 = 0 s = + 2j 53,4 = +2j s-plane 1,2 S = -5 jw (imaginary axis) *x 23 o Creal axis) - 5